OpenStudy (jlyns128):

If f(x) = 2x - 1 + 3 and g(x) = 5x - 9, what is (f - g)(x)?

OpenStudy (eliesaab):

Example h(x)= 5 x -4 k(x)=-2 x +2 (h-k))x)=( 5 x -4) - (2 x +2)=5x-4-2x-2=3x-6 Mimic this to do your poblem

likeabossssssss (likeabossssssss):

^^

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)